OpenStudy (anonymous):

If cot θ = cos θ / sin θ , then tan 2 θ = __________ .

OpenStudy (anonymous):

=sin 2 theta / cos 2 theta. = 2 tan theta/1- (tan theta)^2

OpenStudy (anonymous):

\[\tan 2 \theta=\frac{ 2\tan \theta }{ 1-\tan ^2 \theta }\]